# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v3

@@ 3.0.5 Hybrid Methods
Wu et al. (2024) implemented the hybrid SSL pre-training framework Enhanced SSL with Masked Autoencoders (ESLA). This innovative approach integrates both contrastive learning (CL) and MAE frameworks to enhance SSL. ESLA improves model generalization and representation while addressing issues stemming from competing features during data enhancement. This is achieved through a masking mechanism in the CL framework and by blending representations of different views in the latent space.
@@
Quan et al. (2023) [74] introduced a novel SSL model, based on the global contrast-masked autoencoder (GCMAE), which adeptly extracts both global and local features from pathological images. GCMAE leverages masking image reconstruction and contrastive learning as self-supervised pretext tasks, enabling the encoder to proficiently depict local-global features. This model comprises of an initializer, an encoder, a tile feature extractor, and a global feature extractor facilitating image reconstruction and contrastive learning tasks. In this investigation, an asymmetric encoder-decoder architecture is employed. ViT serves as the backbone for the encoder, pretrained using MAE. Its module related to patches feature extraction is comprised of eight transformer blocks, forming an asymmetric structure in conjunction with an encoder featuring of about 12 transformer blocks (ViT-base). This asymmetry facilitates decoupling of the encoder and decoder, fostering the acquisition of more generalized representations. Extraction of global feature is achieved via utilizing contrastive learning. The hidden representation Vpatch(vis) is used to reconstruct image through decoder and update a memory bank (B). This updation of B is executed to store the global features utilizing momentum coefficient t. The dynamically update is conducted to pursue the feature vectors of image data using memory bank as a fixed size queue. The various features that are randomly picked from the memory bank are used as a negative samples that helps to reduce the effects of the constraint caused by batch size on the heavy performance. Notably, the momentum adjustment mechanism is designed distinctively from MoCo’s model parameter update, it employ an encoder and a memory bank to facilitate contrastive learning. Vpath(vis) is depicted as the encoders output in the current epoch in the latent representation, where as the vpatch(vis) represented the previous epoch latent representation stored in B. Particularly, the latent demonstrations at the output of encoder as vpatch(vis) in the current epoch, The cost function combines both, the weighted sum of MSE loss for tile feature extraction and the NCE loss for global feature extraction. This formulation aims to minimize the gap between similar features, enhancing the model’s generalization and accuracy, particularly in cross-dataset transfer learning tasks.
@@ 4 Regularization Techniques
SimCLR [22] utilizes a straightforward contrastive learning framework, where its regularization is primarily through aggressive data augmentation techniques. These augmentations include random cropping, resizing, color distortion, and Gaussian blur, enabling the model to learn invariant and robust features. This is similar to how techniques like dropout and data augmentation in traditional settings prevent over-reliance on specific input features.
DINO (Self-Distillation with No Labels) [45] utilizes self-distillation for regularization, where the student network mimics the teacher network, updated as an exponential moving average of the student’s parameters. This method is reminiscent of techniques like early stopping and weight decay, which aim to prevent overfitting by smoothing the learning trajectory.
@@
SwAV (Swapping Assignments between Views) [27] regularization is unique, involving clustering approaches and consistency enforcement between cluster assignments for different views of the same image. This approach is akin to enforcing a constraint (regularization) on the feature space to enhance generalizability.
@@ 5.1 Performance Metrics for SSL in ViTs
The performance of SSL methods in ViTs is typically evaluated on downstream tasks, such as image classification, object detection, and semantic segmentation. The performance metrics used to evaluate these tasks include accuracy, precision, recall, F1 score, and mean average precision (mAP). These metrics provide a quantitative measure of the performance of the SSL method in comparison to other methods.
@@ 5.1.1 Accuracy
Accuracy is a measure that represents the true positives (both true negatives and true positives) of the under observation data. Accuracy and error rate are mutually exclusive and range between 0 to 100. Accuracy is computed by dividing the number of correct predictions by the total prediction number as:
@@
5.1.2 Precision
Precision, also known as a measure of quality, quantifies the total number of correct positive predictions. More specifically, it is computed as the number of rightly predicted positive samples divided by the total number of positive examples that were predicted as:
@@ 5.1.2 Precision
Note that the precision is usually preferred in the cases when the false positives are rarely desireable whereas false negatives are tolerable in a learning system.
5.1.3 Recall
Recall is the total number of data samples that a learning network correctly predicts as belonging to the positive class as:
$Recall=\frac{TruePositives}{TruePositives+FalseNegatives}$ (3)
This gives an estimate of the sensitivity of the network. In other words, this provides an idea of the probability that a sample from actual positive class will be classified as positive. It may be noted that recall as a metric is preferable when it is not desired to miss any relevant sample even if it means classifying some samples incorrectly as positive class.
5.1.4 F1 Score
F1 score is computed by taking harmonic mean of Precision and Recall as:
$F1Score=\frac{2\times Precision\times Recall}{Precision+Recall}$ (4)
In simple form, it becomes:
$F1Score=\frac{TruePositives}{TruePositives+\frac{FalsePositives+FalseNegatives}{2}}$ (5)
The reason harmonic mean is used instead of geometric mean is to to balance the values for both Precision and Recall. The more the precision and recall scores deviate from each other, the worse the harmonic mean.
Note that F1 Score is usually preferable when the classes are imabalanced in the dataset which is usually the case in real-world datasets.
@@ 5.3 Loss functions
Machine learning algorithms must include loss functions to gauge how well a model is doing on the job at hand and acts as a feedback signal to adjust the model’s parameters during training. By identifying and quantifying the discrepancy between expected output and the actual target labels, they play a critical part in directing the learning process. Over the past few years, contrastive learning has appeared as an effective approach in the realm of unsupervised representation learning, revolutionizing how neural networks can be trained without labeled data. This approach hinges on the idea of learning meaningful representations by increasing the similarity among positive pairs and decreasing the similarity of negative pairs. Among the several contrastive loss functions, the InfoNCE loss [8] (Noise Contrastive Estimation) stands as a common choice. It operates by introducing noise to the input data to create negative pairs, enabling the model to differentiate between real positive pairs and artificially generated negatives. Another widely used loss is the NT-Xent loss [22] (Normalized Temperature-Scaled Cross Entropy), which extends InfoNCE by incorporating a temperature parameter to scale the pairwise similarities. Through a softmax function, it computes the negative log-likelihood of positive pairs, enhancing the overall contrastive learning process. SimCLR loss (Simple Contrastive Learning) represents yet another approach, harnessing multiple augmentations of the input data to build diverse views of the same instance. It seeks to minimize the similarity between views of different instances while increasing the correspondence between varied augmented views of the same instance. For learning representations or embeddings, the triplet loss [81] has gained popularity. It utilizes an anchor, a positive example that is close to it, and a negative example that is distant. The objective of the loss function is to reduce the distance between the anchor and a positive example while increasing the distance from the negative example, effectively creating a margin between them. The concept of margin loss also focuses on pairwise comparisons, but its objective is to ensure a margin between the representations of positive and negative pairs. By creating pairs of instances and optimizing the model, the loss pushes positive pairings closer together while pushing negative pairings farther apart. In response to the shortcomings of triplet loss, the Circle Loss[82] was recently developed. By establishing a decision boundary in a hypersphere around instances and incorporating a flexible buffer based on the instance distances, it encourages positive pairs to lie within the boundary while keeping negative pairs outside. These are just a few instances of the wide array of contrastive loss functions employed in various research works. The selection of a loss function is determined by factors like the specific problem, the nature of the data, and the required characteristics of the learned representations. Researchers often experiment with multiple loss functions to determine the most suitable one for their specific task [83]. As contrastive learning progresses, more innovative loss formulations and improvements are likely to emerge, opening up new possibilities for unsupervised representation learning .
@@ 6 Comparative Analysis of SSL Mechanisms
In this section, we will provide a comparative analysis of state-of-the-art SSL (SSL) mechanisms employing ViTs. Various SSL methods used different networks as their backbone networks. For a fair comparison, we compare different SSL methods employing ViT-B/16 [10] as their backbone network.
@@ 6.2.1 Object detection and instance segmentation
We present a comparative analysis of SOTA self-supervised approaches for downstream tasks such as object detection and instance segmentation (as shown in Table 5. For a fair comparison, we compare methods having ViT-B [10] as the backbone architecture. The methods are trained over COCO [79] dataset considering Mask R-CNN [98] as the task head for object detection and instance segmentation and evaluated using average precision $AP^{box}$ and $AP^{mask}$, respectively. The results in Table 5 demonstrate that TEC [88], dBOT [89], and MSG-MAE [85] outperform other SOTA SSL methods over object detection and obtain $AP^{box}$ scores of 52.3, 52.7, and 52.3 respectively. On the instance segmentation, MSG-MAE [85], TEC [88], and dBOT [89] show significant promising performance compared to other SSL methods and achieve 48.8, 47.2, and 45.7 $AP^{mask}$ scores, respectively.
@@ 6.2.3 Transfer Classification Comparison
To further study the generalizability of the SSL methods, we compare their transfer learning capabilities in terms of accuracy. To do so, the pre-trained methods are fine-tuned over small datasets such as CIFAR-10 [80], CIFAR-100 [80], iNaturalist18 [102] (iNa18), iNaturalist19 [102] (iNa19), Flowers [103], and Cars [104]. The results in Table 6 Deit-III [92] demonstrate outstanding performance over state-of-the-art SSL methods and achieve 99.3 and 92.5 accuracy over CIFAR-10 and CIFAR-100 datasets, respectively. We also note that dBOT [89] illustrates better performance over CIFAR-10, iNa18, and iNa19 datasets, and show accuracy of 99.3, 77.9, and 81.0, respectively. Similarly, iBOT [86] exhibits the best performance over Flowers and Cars datasets while obtaining 98.9 and 94.3 accuracy, respectively.
@@ 7.1.3 Challenges of Data Augmentation
  • Designing of augmentation strategies: Generally, traditional methods for data augmentation like flipping, random cropping, and color jittering are used. It is not a simple task to design a strategy for effective data augmentation in SSL. Exploring novel and appropriate augmentations which exploit certain characteristics of ViTs is an open challenge [124].

### v4

@@ 3.1.4 Hybrid Methods
Tianhu et al. (2025) [78], proposed a hybrid deep learning model MAEMC-NET, which combines masked autoEncoders for reconstructing masked CT patches and Momentum Contrast (MoCo) for learning discriminative features from different CT views.
Wu et al. (2024) [79] implemented the hybrid SSL pre-training framework Enhanced SSL with Masked Autoencoders (ESLA). This innovative approach integrates both contrastive learning (CL) and MAE frameworks to enhance SSL. ESLA improves model generalization and representation while addressing issues stemming from competing features during data enhancement. This is achieved through a masking mechanism in the CL framework and by blending representations of different views in the latent space.
@@
Quan et al. (2023) [80] introduced a novel SSL model, based on the global contrast-masked autoencoder (GCMAE), which adeptly extracts features from pathological images including both local and global features. GCMAE leverages masking image reconstruction and contrastive learning as self-supervised pretext tasks, enabling the encoder part to proficiently depict local and global features. This model is comprises of an initializer, encoder, a tile feature extractor, and a global feature extractor facilitating image reconstruction and contrastive learning tasks. In this investigation, an asymmetric encoder-decoder architecture is employed. ViT serves as the backbone for the encoder, pretrained using MAE. The module responsible for extracting features from patches consists of 8 transformer blocks, creating an asymmetric configuration when combined with an encoder comprising approximately 12 transformer blocks, as seen in the ViT-base architecture. This asymmetry facilitates decoupling of the encoder and decoder, fostering the acquisition of more generalized representations. Extraction of global feature is achieved via utilizing contrastive learning. The hidden representation Vpatch(vis) is used to reconstruct image through decoder and update a memory bank (B). This updation of B is executed to store the global features utilizing momentum coefficient t. The dynamically update is conducted to pursue the feature vectors of image data using memory bank as a fixed size queue. The features randomly selected from memory bank are utilized as negative samples, that helps to reduce the effects of the constraint caused by batch size on the heavy performance Notably, the momentum adjustment mechanism is designed distinctively from MoCo’s model parameter update, it employ an encoder structure and a memory bank to facilitate contrastive learning. Vpath(vis) is depicted as the encoders output in the current epoch in the latent representation, where as the vpatch(vis) represented the previous epoch latent representation stored in B. Particularly, the latent demonstrations at the output of encoder as vpatch(vis) in the current epoch, The cost function combines both, the weighted sum of MSE loss for tile feature extraction and the NCE loss for global feature extraction. This formulation aims to minimize the gap between similar features, enhancing the model’s generalization and accuracy, particularly in cross-dataset transfer learning tasks.
@@ 4.2 MoCo
4.3 SimCLR
SimCLR [22] utilizes a straightforward contrastive learning framework, where its regularization is primarily through aggressive data augmentation techniques. These augmentations consist of random cropping, resizing, color manipulation, and the application of Gaussian blur, enabling the model to learn invariant and robust features. This is similar to how techniques like dropout and data augmentation in traditional settings prevent over-reliance on specific input features.
4.4 DINO
SINO (Self-Distillation with No Labels) [45] utilizes self-distillation for regularization, where the student network mimics the teacher network, updated as an exponential moving average of the student’s parameters. This method is reminiscent of techniques like early stopping and weight decay, which aim to prevent overfitting by smoothing the learning trajectory.
4.5 Barlow Twins
@@
4.6 SwAV
SwAV (Swapping Assignments between Views) [27] regularization is unique, involving clustering approaches and consistency enforcement between cluster assignments for alternate projections of the single image. This approach is akin to enforcing a constraint (regularization) on the feature space to enhance generalizability.
4.7 TiCo
@@ 5.1.1 Accuracy
Precision, also known as a measure of quality, quantifies the portion of correctly predicted positive instances. Specifically, it is calculated as the ratio of correctly predicted positive samples to the total number of samples predicted as positive:
@@ 5.3 Loss functions
Machine learning algorithms must include loss functions to gauge how well a model is doing on the job at hand and acts as a feedback signal to adjust the model’s parameters during training. By identifying and quantifying the discrepancy between expected output and the actual target labels, they play a critical part in directing the learning process. Over the past few years, contrastive learning has appeared as an effective approach in the realm of unsupervised representation learning, revolutionizing how neural networks can be trained without labeled data. This approach hinges on the idea of learning meaningful representations by increasing the similarity among positive pairs and decreasing the similarity of negative pairs.
5.3.1 InfoNCE Loss
InfoNCE loss (Information Noise Contrastive Estimation) [44] is a loss function commonly used in contrastive learning, particularly in self-supervised learning tasks. It operates by introducing noise into the input data to create negative pairs, enabling the model to differentiate between real positive pairs and artificially generated negatives.
Given a query $q$ and a set of keys ${k_{0},k_{1},\dots,k_{N}}$, where $k_{0}$ is the positive key and $k_{1},\dots,k_{N}$ are negative keys, the InfoNCE loss is formulated as:
$\mathcal{L}=-\mathbb{E}\left[\log\frac{\exp(s(q,k_{0})/\tau)}{\sum_{j=0}^{N}\exp(s(q,k_{j})/\tau)}\right]$ (3)
5.3.2 NT-Xent/SimCLR Loss
NT-Xent (Normalized Temperature Scaled Cross Entropy) [22], which extends InfoNCE by incorporating a temperature parameter to scale the pairwise similarities. Through a softmax function, it computes the negative log-likelihood of positive pairs, enhancing the overall contrastive learning process.
The NT-Xent loss is defined for a pair of augmented views $(i,j)$ of the same instance in a batch. First, the similarity between embeddings $z_{i}$ and $z_{j}$ is computed using cosine similarity:
$sim(z_{i},z_{j})=\frac{z_{i}\cdot z_{j}}{|z_{i}||z_{j}|}$ (4)
The loss function for a positive pair $(i,j)$ is given as:
$\mathcal{L}{i,j}=-\log\frac{\exp(sim(z_{i},z_{j})/\tau)}{\sum\limits{k\neq i}\exp(sim(z_{i},z_{k})/\tau)}$ (5)
5.3.3 Triplet Loss
In many machine learning tasks, particularly in face recognition, person identification, and similarity-based retrieval, a key challenge is learning meaningful feature representations that ensure similar instances are clustered together while dissimilar ones are well separated. Traditional classification-based losses struggle in scenarios with high intraclass variability and low interclass differences, leading to misclassifications. Triplet loss [87] addresses this issue by using a relative distance-based approach, which ensures that an anchor sample is closer to a positive sample (the same class) than to a negative sample (different class) by a specified margin. By simultaneously minimizing the anchor-positive distance and maximizing the anchor-negative distance, Triplet Loss enforces a structured embedding space that enhances discrimination. This approach has proven highly effective in deep metric learning tasks, such as FaceNet, where learning compact and well-separated embeddings is crucial for robust face verification, clustering, and retrieval in unseen data scenarios.
Given an anchor sample $x_{a}$, a positive sample $x_{p}$, and a negative sample $x_{n}$, the triplet loss is formulated as follows:
$\mathcal{L}=\sum_{i}\max\left(d(f(x_{a}^{i}),f(x_{p}^{i}))-d(f(x_{a}^{i}),f(x_{n}^{i}))+\alpha,0\right)$ (6)
5.3.4 Barlow Twins Loss
In self-supervised learning, achieving meaningful feature representations without labeled data is challenging, especially when avoiding collapse, where embeddings become trivial or redundant. Unlike contrastive approaches that rely on distinguishing between different samples, Barlow Twins Loss takes a redundancy reduction approach to ensure diverse and informative features. It aligns representations of augmented views of the same sample, while reducing unnecessary correlations across feature dimensions. This is done by computing the cross-correlation matrix between embeddings from two views and encouraging it to be as close to the identity matrix. By minimizing redundancy without relying on negative pairs, Barlow Twins enables models to learn robust and generalizable representations, making it effective for tasks where contrastive learning may be limited by batch size constraints or reliance on negative samples.
Given two embeddings $Z_{A}$ and $Z_{B}$ corresponding to two different augmentations of the same sample, we define the cross-correlation matrix $C$ as:
$C_{ij}=\frac{\sum_{b}z_{A,b}^{i}z_{B,b}^{j}}{N}$ (7)
where $N$ is the batch size, and $z_{A,b}^{i}$ represents the $i$-th feature of the embedding $Z_{A}$ for the sample $b$.
The Barlow Twins Loss is then formulated as:
$\mathcal{L}=\sum_{i}(1-C_{ii})^{2}+\lambda\sum_{i\neq j}C_{ij}^{2}$ (8)
5.3.5 Reconstruction Loss - MSE - L1 Loss
Reconstruction loss is commonly used in tasks such as autoencoders and generative models to measure how well a reconstructed output matches the original input.
Mean Squared Error (MSE) Loss
MSE computes the squared difference between predicted and actual values:
$\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$ (9)
MSE penalizes large errors more due to the squared term, making it sensitive to outliers.
L1 Loss (Mean Absolute Error, MAE)
L1 loss calculates the absolute difference between predicted and actual values:
$\mathcal{L}_{L1}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$ (10)
Unlike MSE, L1 Loss is more robust to outliers as it does not square the errors. It is often used when preserving small details in reconstructions is important.
5.3.6 Circle Loss
In response to the shortcomings of triplet loss, the Circle Loss[88] was recently developed. By establishing a decision boundary in a hypersphere around instances and incorporating a flexible buffer based on the instance distances, it encourages positive pairs to lie within the boundary while keeping negative pairs outside.
These are just a few instances of the wide array of contrastive loss functions employed in various research works. The selection of a loss function is determined by factors like the specific problem, the nature of the data, and the required characteristics of the learned representations. Researchers often experiment with multiple loss functions to determine the most suitable one for their specific task [89]. As contrastive learning progresses, more innovative loss formulations and improvements are likely to emerge, opening up new possibilities for unsupervised representation learning .
@@ 6 Comparative Analysis of SSL Mechanisms
In this section, we will provide a comparative analysis of state-of-the-art SSL (SSL) mechanisms employing ViTs. Various SSL methods used different networks as their backbone networks. For an unbiased comparison, we compare different SSL methods employing ViT-B/16 [10] as their backbone network.
@@ 6.2.1 Object detection and instance segmentation
We present a comparative analysis of SOTA, a self-supervised approaches for downstream tasks such as object detection and instance segmentation (as shown in Table 5. For an unbiased evaluation, we compare techniques having ViT-B [10] as the backbone architecture. The methods are trained over COCO [85] dataset considering Mask R-CNN [104] as the task head for object detection and instance segmentation and evaluated using average precision $AP^{box}$ and $AP^{mask}$, respectively. The results in Table 5 demonstrate that TEC [94], dBOT [95], and MSG-MAE [91] outperform other SOTA SSL methods over object detection and obtain $AP^{box}$ scores of 52.3, 52.7, and 52.3 respectively. On the instance segmentation, MSG-MAE [91], TEC [94], and dBOT [95] show significant promising performance compared to other SSL methods and achieve 48.8, 47.2, and 45.7 $AP^{mask}$ scores, respectively.
@@ 6.2.3 Transfer Classification Comparison
To further study the generalizability of the SSL methods, we compare their transfer learning capabilities in terms of accuracy. To do so, the pre-trained methods are fine-tuned over small datasets such as CIFAR-10 [86], CIFAR-100 [86], iNaturalist18 [108] (iNa18), iNaturalist19 [108] (iNa19), Flowers [109], and Cars [110]. The results in Table 6 Deit-III [98] demonstrate outstanding performance over state-of-the-art SSL methods and achieve 99.3 and 92.5 accuracy over CIFAR-10 and CIFAR-100 datasets, accordingly. We also note that dBOT [95] illustrates better performance over CIFAR-10, iNa18, and iNa19 datasets, and show accuracy of 99.3, 77.9, and 81.0, respectively. Similarly, iBOT [92] exhibits the best performance over Flowers and Cars datasets while obtaining 98.9 and 94.3 accuracy, respectively.
@@ 7.1.3 Challenges of Data Augmentation
  • Designing of augmentation strategies: Generally, traditional methods for data augmentation techniques includes flipping, random cropping, and color jittering. It is not a simple task to design a strategy for effective data augmentation in SSL. Exploring novel and appropriate augmentations which exploit certain characteristics of ViTs is an open challenge [130].